### v1

# GreekBarBench: A Challenging Benchmark for Free-Text Legal Reasoning and Citations

Abstract
We introduce GreekBarBench, a benchmark that evaluates LLMs on legal questions across five different legal areas from the Greek Bar exams, requiring citations to statutory articles and case facts. To tackle the challenges of free-text evaluation, we propose a three-dimensional scoring system combined with an LLM-as-a-judge approach. We also develop a meta-evaluation benchmark to assess the correlation between LLM-judges and human expert evaluations, revealing that simple, span-based rubrics improve their alignment. Our systematic evaluation of $13$ proprietary and open-weight LLMs shows that even though the best models outperform average expert scores, they fall short of the 95th percentile of experts.
GreekBarBench: A Challenging Benchmark for Free-Text Legal Reasoning and Citations
Odysseas S. Chlapanis1,2Dimitrios Galanis3Nikolaos Aletras4Ion Androutsopoulos1,2 1Department of Informatics, Athens University of Economics and Business, Greece 2Archimedes, Athena Research Center, Greece 3Athena Research Center, Greece 4University of Sheffield, United Kingdom
1 Introduction
As legal AI assistants become increasingly prevalent, the need for realistic legal LLM benchmarks has never been more imperative. [^1] https://www.abajournal.com/web/article/aba-tech-report-finds-that-ai-adoption-is-growing-but-some-are-hesitant Most widely used legal Natural Language Processing (NLP) benchmarks (Chalkidis et al., 2022; Niklaus et al., 2023) focus on classification tasks, e.g., legal judgement prediction (Aletras et al., 2016), which have been criticized (Medvedeva and Mcbride, 2023) for being more constrained and less representative than real-world tasks. Even more recent LLM-focused legal benchmarks (Guha et al., 2023; Fei et al., 2024; Joshi et al., 2024) do not go beyond closed-form questions (e.g., multiple-choice questions), failing to capture the true complexity of legal reasoning in practice, which involves identifying, analyzing and synthesizing relevant information to reach a conclusion. Unfortunately, most existing benchmarks with challenging legal questions and free-text responses are proprietary and thus inaccessible to the research community. [^2] https://www.vals.ai/benchmarks
Facts
$[1]$ Antonis visited his dermatologist, Ioannis, to remove facial skin tags.
$[2]$ Ioannis’s assistant, Penelope, accidentally applied pure acetic acid due to a mislabeled bottle, causing burns on Antonis’s face.
$[3]$ He needed plastic surgery costing €2,500 and is now seeking these costs plus €75,000 for moral damages.
Question
Which individuals are liable for the injury?
Relevant Legal Context
Civil Code 914: Anyone who unlawfully causes damage must compensate the victim.
Civil Code 922: An employer is liable for unlawful damages caused by their employee during work.
Ground Truth Answer
Ioannis is responsible vicariously for Penelope’s
actions $[2]$(Civil Code 922)and Penelope is directly
liable for her negligence $[2]$(Civil Code 914).Hence,
both are liable and must compensate Antonis.
Table 1: Cropped example (English translation) from GreekBarBench. The answer requires multi-hop reasoning and citing legal articles and case facts. The spans corresponding to the scoring dimensions are highlighted in color: Facts (green), Cited Articles (blue) and Analysis (orange). Important spans are marked in bold and cited facts are denoted by square brackets. The complete example is presented in Appendix C.
Another challenge is that realistic benchmarks often require costly manual evaluation by legal experts, which limits scalability (Magesh et al., 2025; Martin et al., 2024). Automatic evaluation, using the LLM-as-a-judge framework (Zheng et al., 2023), is a promising alternative; however, its reliability has not been extensively assessed in legal reasoning (Bhambhoria et al., 2024; Li et al., 2025).
To address these issues, we present the GreekBarBench, a benchmark that evaluates the reasoning capabilities of LLMs on challenging legal questions across five legal areas. The questions are taken from the Greek Bar exams and require open-ended answers with citations to statutory articles and case facts. In addition, we introduce an accompanying benchmark for LLM-judges, designed to measure how well their scores correlate with those of human experts. GreekBarBench is the only Greek dataset for legal reasoning. Our main contributions are the following:
  • GreekBarBench: a challenging legal reasoning benchmark that requires free-text answers citing case facts and statutory articles.
  • GBB-JME: an accompanying dataset with human-evaluated answers from five different LLMs, to assess the quality of candidate LLM-judges in GreekBarBench.
  • A three-dimensional scoring system and an LLM-judge framework based on span-rubrics per dimension (Facts, Cited Articles, Analysis), which aligns well with human expert evaluation.
  • A systematic evaluation of 13 frontier and open-weight LLMs on GreekBarBench, using the best LLM-judge at GBB-JME. Top models surpass average expert performance, but not the 95th percentile of experts.
All resources including the two benchmarks (except for a small semi-private test set) and the prompts are publicly available. [^3] The URL of the GitHub repository will be provided in the camera-ready version.
2 GreekBarBench (GBB)
2.1 Greek Bar Exams
Law graduates in Greece must pass the Greek Bar exam to become licensed attorneys. The exam evaluates candidates through practical legal questions across five key areas of law: Civil Law, Criminal Law, Commercial Law, Public Law, Lawyers’ Code. Greece’s legal system is statutory, meaning that laws are derived from legal code documents (statutes), rather than from judicial precedents (case law). The exams are open-book; candidates have access to legal code documents and are expected to cite statutory articles from them in their answers.
The available documents include: the Civil Code and Civil Procedure Code, the Criminal Code and Criminal Procedure Code, eight Commercial Law codes, eleven Public Law codes, as well as the Lawyers’ Code and the Code of Ethics for Legal Practice (see Table 3). Candidate lawyers typically approach the exam by first studying the case facts to identify the relevant legal issues. They then navigate the legal code documents to find the relevant chapter and pinpoint the exact statutory article within it to cite in support of their arguments.
2.2 Task
Benchmark Lang Citations Multi-Hop Free-Text Judge Eval
LegalBench en
LexEval (5.4) zh
CaseGen zh
OAB-Bench por
LLeQA fr
GBB (Ours) el
Table 2: Comparison of legal benchmarks. GreekBarBench uniquely encompasses all challenging features essential for evaluation in realistic and practical scenarios. ‘Lang’: language of dataset. ‘Citations’: legal articles. ‘Multi-hop’: reasoning using multiple sources. ‘Free-text’: open-ended responses. ‘Judge Eval’: manual evaluations to compare LLM-judges.
Each instance in GreekBarBench is taken from Greek Bar exam papers. The input consists of (1) the case facts, (2) the legal question, and (3) a collection of potentially relevant chapters of statutory articles. The desired output is the correct free-text answer to a legal question, providing an analysis with citations to the case facts and the applicable statutory articles. The primary challenges are to discern significant facts, to identify the applicable articles and, finally, to analyze the outcomes of the application of the articles to answer the question (see the example in Table 1). These attributes make GreekBarBench unique compared to other legal benchmarks (Table 2).
2.3 Dataset Statistics
We collect a total of 65 exam papers; 13 exam papers from each of the five aforementioned areas. The papers and suggested solutions are publicly available in a booklet in PDF format, spanning from 2015 to 2024.444The booklet is available at https://www.lawspot.gr/nomika-nea/panellinios-diagonismos-ypopsifion-dikigoron-themata-exetaseon-kai-endeiktikes-3. The authors permit its distribution for academic research only. The booklet is converted to text format and further processed (§ 2.5) to prepare the dataset. Each exam paper includes 4.7 questions on average, resulting in a total of 310 samples. We keep the questions from 2024 (22 in total) as a semi-private test set, to avoid data contamination. [^5] With ‘semi-private’ we mean that the test set is not public, but the raw data sources are available. The remaining 288 samples comprise the public test set. The semi-private set will be updated each year with two more exam papers from each legal area (there are two examinations per year), and made publicly available.
Answering exam questions requires citing articles from 25 legal code documents, which we collect from the same source website as the exams.666www.lawspot.gr Detailed statistics for these documents are presented in Table 3 per legal area. Articles are cited 931 times in total, across all exam questions. The articles within each legal code document are grouped thematically into chapters. The total number of citable articles is more than 16 thousand.
Law Areas Samples Legal Total Cited Context
Codes Articles Articles (tokens)
Civil 71 2 3,264 286 87k
Criminal 53 2 1,253 186 58k
Commercial 58 8 4,177 159 29k
Public 71 11 2,912 118 67k
Lawyers 57 2 4,476 182 66k
Total 310 25 16,082 931 62k
Table 3: Summary of dataset statistics. ‘Legal Codes’ indicates the number of distinct legal code documents in each area. ‘Cited Articles’ is the total number of citations to legal code articles. ‘Context’ denotes the average token count of the relevant legal context (chapters of legal code) provided in the input of candidate LLMs.
2.4 Relevant Legal Context
As mentioned in Section 2.1, the Greek Bar exams are open-book, allowing candidate lawyers to navigate legal code documents to identify relevant statutory articles for the presented case. Simulating this setup presents several challenges. One approach would be implementing a standard Retrieval-Augmented Generation (RAG) pipeline, using sparse (e.g., BM25) or dense retrievers (Karpukhin et al., 2020) to select the $k$ most ‘relevant’ articles for inclusion in the LLM’s input. However, this approach suffers from three significant limitations: a) candidate lawyers taking the exams do not have access to such retrieval tools, making direct comparisons with human performance problematic; b) retrievers are prone to errors, creating a substantial risk that even with large values of $k$, the ground truth articles might not appear among the top retrieved articles; and c) as demonstrated by Krishna et al. (2025), benchmarking RAG systems requires testing multiple configurations with varying values of $k$ and, ideally, different retriever models, complicating fast integration of new LLMs.
Instead, we adopt a simplified yet challenging approach. For each legal case, we automatically collect all the ground truth articles cited across all questions pertaining to that case (using regular expressions), and identify the chapters containing the articles. We hypothesize that all the articles of the identified chapters have substantial relevance for all questions of the particular case. Therefore, we provide the entire text (all articles) of the identified chapters as the relevant legal context in the LLM input for every question of the particular case. Hence, for each question, the input context contains all the ground truth articles of the question, along with all the other articles from their chapters, as particularly close distractors. It also includes additional distractor articles from other chapters related to other questions of the same case, which increase the difficulty of citing the correct articles.
The total length of the legal context fed to the LLM per question is 62k tokens on average (Table 3). This makes the task manageable for recent LLMs with context windows exceeding 100k tokens, while still presenting a significant challenge, as LLMs often struggle to extract crucial information from extended contexts (Liu et al., 2024).
2.5 Fact Segmentation
To enhance the evaluation process (both manual and automatic), we require citations to facts in candidate answers, though this is not mandatory in official exams. Fact-citations help legal professionals in practical applications by enabling more efficient verification of answer accuracy. To facilitate fact-citation, we present the case facts as a numbered list of sentences (as shown in Table 1). For segmentation, we employ the Segment-Any-Text neural model (Frohmann et al., 2024).
2.6 Three-Dimensional Scoring System
The official evaluation committee of the Greek Bar Exams grades candidate answers on a scale of 1 to 10. Without explicit guidelines, they grade by comparing answers to the ground truth based on their discretion. Drawing inspiration from established legal research and evaluation practices (Clark and DeSanctis, 2013), and guided by our legal expert annotators (§ 5.2), we develop a novel three-dimensional scoring system to improve the evaluation process for the benchmark. The proposed approach assesses legal reasoning across three dimensions: the Facts, the Cited Articles, and the Analysis. Each dimension is rated on a scale of 1 to 10, and the final score is the average of these three. This system allows explainability through detection of specific shortcomings in the reasoning abilities of LLMs. The Facts score measures understanding of case facts; the Cited Articles evaluates the accuracy and interpretation of cited legal articles; and the Analysis evaluates the ability to apply legal articles to the facts and reach conclusions. For instance, a low Facts score indicates hallucinations, a low Cited Articles score shows difficulty in identifying applicable articles, and a low Analysis score reveals weakness in legal reasoning.
3 Automatic Evaluation
To address the evaluation of free-text answers without the prohibitive cost of manual annotations, we use the LLM-as-a-judge framework (Zheng et al., 2023). LLM-judges can be categorized into two primary types: (a) pairwise LLM-judges, which evaluate two candidate answers and determine which is preferred (or declare a tie), and (b) grading LLM-judges, which assign an integer score to each individual candidate answer (Zheng et al., 2023). In our work, we focus on grading LLM-judges to allow cost-effective integration of new participant LLMs without the overhead of quadratically increasing pairwise comparisons.
To improve the alignment of LLM-judges with human expert annotators, we propose novel span-based rubrics; i.e., evaluator instructions in the form of annotated spans per question. These spans will guide the LLM-Judge in what to assess in the candidate answers. However, even with these question-specific rubrics, replicating the nuanced evaluation of human experts, especially in complex tasks like legal writing, cannot be guaranteed. For this reason, we also include a framework to meta-evaluate whether LLM-judges are suitable proxies for human evaluation on GreekBarBench.
3.1 Simple LLM-Judge
As an initial approach, we designed a straightforward prompt for a simple LLM-judge. The prompt outlines the evaluation task and explicitly defines the criteria for the Facts, Cited Articles, and Analysis scores. All necessary contextual information is provided; the facts of the case, the specific legal question, the ground truth answer with the cited articles and the candidate answer to be evaluated. This context mirrors the information provided to the human annotators for the manual evaluations (§ 5.2). The required output format is clearly specified: the model must provide an explanation for each score, followed by the integer score. The complete prompt is presented in Appendix B (Fig. 5).
3.2 Span LLM-Judge
According to Clark and DeSanctis (2013), rubrics, i.e., instructions that break down an assignment in identifiable components, can significantly improve the consistency of legal writing evaluation. To construct rubrics for our benchmark, our legal expert annotators marked reference spans (colored text in Figure 1) in the ground-truth answer for each score (Facts, Cited Articles, Analysis). These spans contain the information pertinent to their respective dimension. Each span is then annotated with important span subsets (usually a few words within the span) that are crucial for an answer to be considered correct (bold text in Table 1). Missing this crucial information should result in a lower score. We opted not to assign specific point values or ‘costs’ to each important subset, as previous work with rubrics (Starace et al., 2025; Pires et al., 2025), to minimize the annotation burden. The process involves simply highlighting the three scoring dimensions using different colors and then marking the important subsets within those highlighted sections. The LLM-judge is instructed to determine whether the candidate answer covers the information in the spans and then use this assessment to evaluate each scoring dimension. The complete prompt is presented in Appendix B (Fig. 6).
4 Meta Evaluation
Meta-evaluation of grading LLM-judges aims to quantify the alignment between LLM-generated scores and human expert annotations. Previous research has predominantly relied on Pearson’s or Spearman’s correlation coefficients as primary meta-metrics (Bavaresco et al., 2024; Niklaus et al., 2025), often without substantial justification. However, advancements in meta-evaluation have emerged from the machine translation domain, particularly through the WMT Metrics Shared Task (Freitag et al., 2024, 2023), where automatic evaluation frameworks have been systematically compared and refined. The task aims to identify optimal metrics for translation quality assessment by comparing system outputs against references. Recent findings demonstrate that state-of-the-art metrics are increasingly LLM-based. The task has revealed that Pearson’s correlation coefficient exhibits vulnerability to outliers, while Spearman’s $\rho$ disregards the magnitude of ranking errors, applying uniform penalties. To address these limitations, WMT has adopted Soft Pairwise Accuracy (SPA) (Thompson et al., 2024), a metric that assigns partial credit for nearly correct rankings, thereby providing an evaluation framework that better reflects the alignment of metrics with human experts.
4.1 Soft Pairwise Accuracy (SPA)
SPA measures the degree of alignment in evaluation confidence between human experts and LLM-judges (or any other automatic metric). For example, if a human expert is confident that one system (e.g., a candidate LLM from GreekBarBench) outperforms another, but the LLM-judge is uncertain, SPA penalizes the judge—even if the ranking happened to be correct. To do this, SPA approximates the confidence of each judge (human or LLM) on each pairwise comparison between systems using p-values of appropriate permutation tests (Fisher, 1935), as detailed below. We use the original implementation. [^7] https://github.com/google-research/mt-metrics-eval Formally, SPA between a metric $m$ and human experts $h$ is defined as:
$\textit{SPA}(m,h)={\binom{N}{2}}^{-1}\sum_{i=0}^{N-1}\sum_{j=i+1}^{N-1}\left(1-\left|p_{ij}^{h}-p_{ij}^{m}\right|\right)$
where $N$ is the number of systems being evaluated, $p_{ij}^{h}$ is the p-value for the hypothesis that system $i$ is better than system $j$ according to human scores, and $p_{ij}^{m}$ is the corresponding p-value according to the metric under evaluation. The term ${\binom{N}{2}}^{-1}$ normalizes the summation by the total number of systems under comparison.
SPA permutation tests:
To estimate confidence of an evaluator (either human or automatic) in a pairwise system comparison, SPA uses permutation tests to calculate the expected mean difference under the null hypothesis that the systems are of equal quality. Specifically, a number of mock systems (1,000 in our experiments, following the original paper) are constructed as follows: for each question in the benchmark, the mock system is assigned either the score of system $i$ or system $j$ at random. The p-value is then computed as the proportion of mock systems for which the differences are greater than or equal to the mean difference between systems $i$ and $j$, as scored by the evaluator.
Figure 1: Manual evaluation by legal expert annotators on the semi-private test set of the 2024 exams.
5 Experiments
5.1 Models
Our experiments evaluate a diverse range of LLMs, comprising proprietary models (from OpenAI, Google, Anthropic) and open-weight models; Deepseek-R1 (DeepSeek-AI et al., 2025), Gemma-3 (Team et al., 2025), and Llama-Krikri-8B [^8] https://huggingface.co/ilsp/Llama-Krikri-8B-Instruct, a model specifically pretrained for the Greek language. We accessed proprietary models and the large open-weight Deepseek-R1 through Application Programming Interfaces (APIs) provided by OpenAI, Google, and AWS. The remaining open-weight models were deployed on a cluster of eight A100 GPUs using the vLLM framework (Kwon et al., 2023). Due to limited resources, we only evaluate a single run for each model. We used the default parameter configurations as specified by each model’s provider.
Generation prompt:
To generate responses from LLM candidates, we designed a system and user prompt for the questions of the benchmark. The system prompt instructs the LLM to answer with citations to Greek statutory articles. The user prompt is structured to first describe the overall task, including clear instructions on the expected output format. Then it provides the numbered facts of the case, the question and the relevant legal context. The original prompts are available in Appendix B.
5.2 Manual Evaluation by Legal Experts
Model Simple-J Span-J Cost
(SPA) (SPA)
[…]
L-Krikri-8B 0.747 0.751 -
Gemma-3-27B 0.819 0.749 -
[…]
In this section we present the manual evaluations that we collected for GBB-JME (§ 5.3), our Judge Meta-Evaluation benchmark for assessing LLM-judges on GreekBarBench. We obtain ground truth evaluations (Facts, Cited Articles, Analysis scores on a scale of 1 to 10) from two expert legal annotators—licensed Greek lawyers with law degrees and practical experience. The annotators were compensated for their time and expertise. They evaluated five LLMs on 87 questions drawn from three exam sessions (2024-A, 2023-A, and 2023-B), resulting in a total of 1,305 annotated samples. The models evaluated on all three exams were Claude-3.7-Sonnet, OpenAI-o1, GPT-4o, and Gemini-2.0-Flash. For the 2024 exam, we included the open-source Llama-3.1-70B; however, due to its poor performance and generation failures on several questions, we replaced it with Deepseek-R1 for the 2023-A and 2023-B exams. Annotations were managed with the open-source platform doccano. [^9] https://doccano.prio.org
The average Krippendorff’s $\alpha$ (Krippendorff, 2011) between the two annotators on the three-dimensional scores was 0.74, and the SPA was 0.85, both indicating a substantial level of inter-annotator agreement (Artstein and Poesio, 2008). For the SPA calculation, we treated one annotator’s scores as ground truth and compared the other annotator’s scores against them. This differs from Section 5.3, where SPA measures the correlation between LLM-generated scores and the aggregated scores of human annotators.
The results for the 2024 exam are shown in Figure 1. All LLMs except Llama-3.1-70B pass the exam. However, all LLMs lag behind the average human expert performance and the 95th percentile (‘top-5%’). The best-performing models are Claude-3.7-Sonnet (7.55) and OpenAI-o1 (7.52).
5.3 Judge Meta-Evaluation (GBB-JME)
We evaluate seven LLMs as judges on our benchmark for meta-evaluation, GBB-JME, using both the Simple-Judge prompt (§3.1) and the Span-Judge prompt with span-based rubrics (§3.2). The results are presented in Table 4. Generally, leading models (GPT-4.1, Gemini-2.0-Flash and GPT-4.1-mini) showed significant improvement when utilizing the Span-Judge prompt, while weaker models (Llama-Krikri-8B, Gemma-3-27B, Gemini-2.0-Lite-Flash, GPT-4.1-nano) struggled with its complexity. GPT-4.1-mini and GPT-4.1 achieved the best performance at 0.856 SPA, and 0.855 SPA, respectively. The competitive performance of the smaller model, aligns with observations reported elsewhere (Niklaus et al., 2025).
Figure 2: Comparison of closed and open-weight LLMs on GreekBarBench with GPT-4.1-mini Span-Judge.
Due to its strong performance and lower cost, we adopted GPT-4.1-mini as the judge for all subsequent evaluations. The total cost for all evaluations using GPT-4.1-mini was under \$60. The open-weight Gemma-3-27B model paired with the Simple-Judge prompt serves as a cost-effective alternative, but we also encourage researchers to evaluate new LLMs as judges on the publicly available GBB-JME benchmark.
Models Civil Law Criminal Law Commercial Law Public Law Lawyers’ Code Overall
f c a avg f c a avg f c a avg f c a avg f c a avg f c a avg
Top-5% - - - 9.00 - - - 10.00 - - - 10.00 - - - 9.20 - - - 9.18 - - - 8.87
Experts - - - 6.80 - - - 8.29 - - - 8.70 - - - 7.70 - - - 7.39 - - - 7.78
Gemini-2.5-F 8.8 8.4 8.4 8.53 8.4 7.9 8.5 8.28 8.6 8.0 8.3 8.27 8.7 8.0 8.1 8.28 8.9 8.5 8.5 8.62 8.7 8.2 8.4 8.40
GPT-4.1 8.8 8.1 8.4 8.44 8.5 8.2 8.2 8.28 8.4 8.1 8.3 8.27 8.6 7.7 8.1 8.14 8.8 8.2 8.4 8.48 8.6 8.0 8.3 8.32
Claude-3-7 8.5 7.2 7.4 7.72 8.2 6.9 7.0 7.37 7.6 6.9 7.4 7.31 8.6 7.2 7.6 7.79 8.5 8.2 8.2 8.29 8.3 7.3 7.5 7.71
GPT-4.1-mini 8.3 7.1 7.3 7.57 7.7 6.4 6.9 7.01 8.4 7.4 7.4 7.76 8.4 7.3 7.6 7.75 8.5 7.6 7.9 7.98 8.3 7.2 7.4 7.63
Gemma-3-27B 7.8 5.5 5.9 6.39 6.7 4.9 4.9 5.51 6.8 5.2 5.6 5.88 8.1 5.9 6.1 6.68 7.8 6.6 6.6 7.01 7.5 5.7 5.8 6.33
L-Krikri-8B 7.0 5.3 5.6 5.95 7.1 5.3 5.1 5.84 6.7 5.3 5.4 5.79 7.9 6.3 6.2 6.78 7.4 6.4 6.4 6.74 7.2 5.7 5.8 6.24
Table 5: Fine-grained comparison of proprietary and small open-weight LLMs on different legal areas: ‘Civil’, ‘Criminal’, ‘Commercial’, ‘Public’, ‘Lawyers’; and for different scoring dimensions: ‘Facts’ (f), ‘Cited Articles’ (c), ‘Analysis’ (a). ‘Experts’/‘Top-5%’ is the average/95th percentile score of candidate lawyers. Best LLM scores are in bold, failed scores are highlighted in red, and scores outperforming experts are highlighted in green.
5.4 Results on GreekBarBench
We conduct an extensive automatic evaluation of $13$ LLMs on GreekBarBench (Figure 2). We use GPT-4.1-mini as the LLM-Judge, employing the ‘Span LLM-Judge’ prompt (§ 3.2). The evaluation includes proprietary models such as GPT-4o, the GPT-4.1 family (GPT-4.1-mini, GPT-4.1-nano), Gemini-2.0-flash, and Claude-3.7-Sonnet (with reasoning disabled), along with the reasoning models OpenAI-o1 and Gemini-2.5-Flash. The open-weight models include the Gemma-3 family(Gemma-3-27B, Gemma-3-12B, and Gemma-3-4B), the specialized Greek model Llama-Krikri-8B-Instruct (Krikri-8B), and the reasoning model DeepSeek-R1.
The experimental results (Figure 2) reveal that Gemini-2.5-Flash (8.4) and GPT-4.1 (8.32) demonstrate the strongest performance on GreekBarBench. They surpass the average legal expert score (7.78), though they still fall short of the 95th percentile (top-5%) of experts (8.87). OpenAI-o1 (7.77) and Claude-3.7-Sonnet (7.71), perform comparably to the average human expert (7.78). The fact that Gemini-2.5-Flash and OpenAI-o1, two reasoning models, are among the top performers, shows that leveraging inference-time reasoning is a key factor for this benchmark. The smallest models, GPT-4.1-nano (5.51) and Gemma-3-4B (4.41) are the only models that fail the exams (passing score: 6.00). The 8B Krikri model surpasses Gemma-12B and achieves performance comparable to the significantly larger Gemma-27B, highlighting the benefit of language-specific pretraining.
Table 5 presents a fine-grained comparison for four of the best proprietary LLMs (Gemini-2.5-Flash, GPT-4.1, Claude-3.7-Sonnet , GPT-4.1-mini) and two small open-weight LLMs (Gemma-3-27B, L-Krikri-8B), reporting their scores across five different legal areas (Civil, Criminal, Commercial, Public, Lawyers) and three scoring dimensions (Facts, Cited Articles, Analysis).
The fine-grained comparison shows that LLMs exhibit consistent performance across all legal areas and scoring dimensions. Notably, in the areas of ‘Civil Law‘, ‘Public Law‘, and ‘Lawyers’ Code‘, four of the LLMs achieve scores surpassing the average human expert performance highlighted in green in Table 5. In contrast, the smaller open models, Gemma-3-27B and L-Krikri-8B (indicated by red), struggle in certain areas, failing to meet the passing grade threshold of 6.00. The second-best model, GPT-4.1, matches the top performer, Gemini-2.5-Flash, in ‘Criminal’ and ‘Commercial Law’, but Gemini-2.5-Flash achieves slightly higher scores in the remaining three areas. Interestingly, even though ‘Civil Law‘ seems to present the greatest challenge for human experts (average score of 6.80), this is not the case for LLMs.
Analyzing performance by dimension provides valuable insights into model capabilities. A key finding is that all models struggle most with the ‘Cited Articles‘ dimension. The higher overall scores of Gemini-2.5-Flash and GPT-4.1 are largely attributable to their stronger performance in the Cited Articles and Analysis dimensions, when compared to Claude-3.7-Sonnet and GPT-4.1-mini. Furthermore, Gemma-3-27B’s stronger performance compared to L-Krikri-8B is primarily attributable to its higher scores in the Facts dimension, while they get identical scores in the other dimensions.
6 Related Work
Legal domain:
In the legal domain, LexGLUE (Chalkidis et al., 2022) and LEXTREME (Niklaus et al., 2023) are established benchmarks for legal classification tasks. LegalBench (Guha et al., 2023) is the standard for evaluating LLMs on legal reasoning via multiple-choice questions. More closely related to our work, task 5.4 of LexEval (Li et al., 2024) uses a similar legal examination for Chinese instead of Greek, but, unlike our approach, LexEval does not provide citations or use LLM-as-a-judge, instead evaluating with the less reliable, overlap-based ROUGE metric (Cohan and Goharian, 2016). LLeQA (Louis et al., 2024) collects pairs of everyday legal questions and answers, including citations to French statutory law articles, but they evaluate using the METEOR metric without measuring its correlation with human experts. CaseGen (Li et al., 2025) on the other hand, assesses document drafting and legal judgment generation in Chinese using the LLM-as-a-judge approach. While they do measure agreement between human and LLM evaluations, they do not compare different prompts or models. Concurrent work, OAB-Bench (Pires et al., 2025), uses data from the Brazilian Bar Examination and also provides the official guidelines as rubrics for the LLM-judges. They do evaluate different LLM-judges, but they only do it for three samples and they do not provide citations to statutory articles. Notably, the complexity of their rubrics necessitates the use of the expensive OpenAI-o1 model for evaluation, thus significantly increasing the overall cost, amounting to approximately \$50 for each LLM evaluated.
LLM-as-a-judge:
LLM-as-a-judge was introduced by Zheng et al. (2023), who meta-evaluated its performance against human preferences for multi-turn chat assistant dialogues. A comprehensive overview of LLM-as-a-judge and meta-evaluation resources can be found in the survey by Gu et al. (2024). Taking this concept further, JudgeBench (Bavaresco et al., 2024) introduced a general-purpose benchmark specifically for the meta-evaluation of LLM-judges. In line with our approach, other studies similarly develop separate benchmarks to meta-evaluate judges on specific tasks (Starace et al., 2025; Niklaus et al., 2025).
Evaluation Rubrics:
Legal research has for long focused on creating rubrics for consistent (human) evaluation of legal writing (Clark and DeSanctis, 2013). The Brazilian Bar exams have made their rubrics for human evaluation available, so the aforementioned OAB-Bench (Pires et al., 2025) provides them to their LLM-judges. Their rubrics consist of a manually annotated ground truth answer with comments and a table with score distributions for each element of the answer. A proprietary benchmark, BigLawBench [^10] https://www.harvey.ai/blog/introducing-biglaw-bench, describes a scoring system that uses two dimensions: the ‘source’ and ‘answer’ scores, which are analogous to our Cited Articles and Analysis. They rely on detailed instructions per question that specify explicitly the attributes that would contribute positively and negatively to the final score of candidate answers. Constructing from scratch either of these approaches is prohibitively expensive, in contrast to our simple, span-based rubrics that only require minimal annotation effort.
Greek NLP:
Important Natural Language Processing resources for the Greek language include classification models (Koutsikakis et al., 2020; Saketos et al., 2024), alongside more recent LLMs pretrained on Greek like Meltemi [^11] Meltemi was excluded from our experiments, because of its relatively small context length of 8 billion tokens. (Voukoutis et al., 2024) and Llama-Krikri [^12] https://huggingface.co/ilsp/Llama-Krikri-8B-Instruct, which we tested in our experiments (§ 5.4). Existing Greek legal datasets cover only classification and summarization tasks (Angelidis et al., 2018; Papaloukas et al., 2021; Koniaris et al., 2023). Although Greek LLM benchmarks exist for other domains, such as finance (Peng et al., 2025) and medicine (Voukoutis et al., 2024), the legal domain currently lacks one.
7 Conclusions
In this work, we introduced GreekBarBench, a benchmark evaluating LLMs on legal questions requiring citations to statutory articles and case facts. We use a comprehensive scoring system and an LLM-judge for automatic evaluation. To ensure judge alignment with human experts, we developed an accompanying meta-evaluation benchmark (GBB-JME) using Soft-Pairwise Accuracy as the meta-metric. The results show that our span-based rubrics specifically designed for this benchmark improve the LLM-judges. The extensive evaluation of 13 LLMs on GreekBarBench revealed that Gemini-2.5-Flash and GPT-4.1 achieved the best performance, surpassing the typical human expert, but also highlighted areas for future improvement.
Limitations
Our benchmark, GreekBarBench, assumes the availability of the relevant legal code chapters for the Relevant Legal Context component (§ 2.4). We did not evaluate the performance of retrieval models on this task, which is a critical step in real-world legal applications and could pose a significant challenge not addressed by our current setup.
A notable limitation is the cost associated with evaluating models using our framework due to the primary LLM-judge being a proprietary model (GPT-4.1-mini). To mitigate this cost, we suggest utilizing Simple-Judge with the open-weight model Gemma-3-27B. While no currently available open-weight model achieves meta-evaluation performance (SPA scores on GBB-JME) on par with GPT-4.1-mini, our public release of the benchmark and meta-evaluation dataset will allow future research to test and use more accurate and cost-effective LLM-judges.
Finally, the reported legal expert performance figures (average and 95th percentile) in our comparisons (§ 1, § 2) should be interpreted as illustrative baselines rather than rigorous head-to-head comparisons under identical evaluation conditions. This is due to inherent limitations in the available human data: detailed participant statistics are not available across all exam papers, and critically, the human scores were determined by the official Greek Bar Examination grading committee, not by our developed LLM-judge framework. Nevertheless, we believe these figures provide valuable intuition regarding the current performance gap between state-of-the-art LLMs and candidate lawyers.
Ethical Considerations
The development and application of legal NLP benchmarks carry significant ethical implications and potential societal impact, particularly concerning fairness, access to justice, and responsible automation (Tsarapatsanis and Aletras, 2021). Therefore, careful consideration of their design and potential uses is essential.
Our research contributes to the development of tools that could potentially assist various types of users, including legal professionals (such as judges and lawyers), students, and individuals seeking to understand legal concepts. It is crucial to emphasize that performance on this benchmark, or any similar research benchmark, should never be considered sufficient justification for deploying automated systems that substitute human experts. We strongly caution against the uncritical reliance on models evaluated solely on benchmark performance for automating legal tasks, making legal decisions, or providing legal advice.
Despite our efforts to make GreekBarBench realistic, as a research benchmark, it overlooks two critical aspects for the safe and reliable deployment of legal AI applications in practice:
  • Data Realism: Real-world legal problems are far more complex and nuanced than the structured, often simplified scenarios found in exam questions (Medvedeva and Mcbride, 2023). They often demand significant legal interpretation, ethical judgment and persuasion, particularly when the law does not provide an explicit answer for a given situation.
  • Safety: Real-world applications must ensure that the AI system handles adversarial attacks effectively. Issues like guiding the decisions of the LLMs with malicious prompting (e.g., jailbreaking), and providing confident, incorrect information when asked legally unanswerable queries are unacceptable (see discussions on AI safety principles  [^13] https://www.anthropic.com/news/core-views-on-ai-safety).
Furthermore, the primary ethical purpose of this work is not to provide a system ready for deployment, but to advance the state of legal NLP evaluation itself. By developing a benchmark that requires free-text generation, incorporates a multi-dimensional scoring system, and uses LLM-judges with explicit evaluation criteria, we aim to encourage the development of more transparent and explainable legal AI models. These features provide greater insight into how models arrive at their answers, moving beyond simple classification or multiple-choice and offering components of explainability which are crucial for gaining trust in AI applications (Medvedeva and Mcbride, 2023).
As already mentioned (§ 2.1), the authors of the solutions of the exam papers have given approval for the public reproduction of this work, with respect to the original and strictly for academic research use. Our ground truth answers are based on the year that each exam was published. This means that if the relevant laws changed in the meantime, the solutions are no longer valid. All cases in the Greek Bar exams are fictional, created solely for educational purposes, and bear no relation to real individuals or actual legal cases.
Acknowledgments
We are grateful to our legal expert annotators, Nasia Makridou and Irene Vlachou, for their diligent work, expertise, and insightful discussions, which were invaluable to this project.
This work was partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. AWS resources were provided by the National Infrastructures for Research and Technology (GRNET), with support from the EU Recovery and Resilience Facility.
Appendix A Annotator Instructions
In this section we present the instructions given to the two legal expert annotators. The annotators possessed prior experience with the evaluation task, having previously taken the exams themselves. This existing expertise allowed for concise instructions. For the general evaluation of LLM-generated answers for the manual evaluation (§ 5.2), the instruction (translated to English) was simply to:
“Evaluate the candidate answers on each scoring dimension (Facts, Cited Articles, and Analysis).”
For the creation of text spans for Span-Judge (§ 3.2), annotators were instructed to:
“Highlight the text-spans that correspond to each scoring dimension (Facts, Cited Articles, and Analysis). Highlight the most important subsets of these spans with the label important.”
Appendix B Complete Prompts
In this section we present the complete system (Fig.3) and user (Fig.4) prompts given to candidate LLMs for generation of answers, as well as the complete system prompts given to LLM-judges for the Simple- (Fig.5) and Span-Judge (Fig.6).
Figure 3: System prompt for generation given to candidate LLMs.
Figure 4: User prompt for generation given to candidate LLMs.
Figure 5: System prompt given to Simple-Judge LLM-judges.
Figure 6: System prompt given to Span-Judge LLM-judges.
Appendix C Complete Dataset Example
In this section we present the complete version of the example that we presented in Table 1. We show the complete Facts and Question (Fig. 7), the Relevant Legal Context (Figures 8 and 9), the complete Ground Truth Answer (Fig. 10), the candidate answer by Gemini-2.5-Flash (Fig. 11) and evaluations of the candidate answer by the legal experts and the LLM-judge (Fig.12).
Figure 7: Complete Facts and Question (original and below translated in English), as given in to the candidate LLMs, for the example in Table 1.
Figure 8: The Chapters of the Relevant Legislation context given to candidate LLMs, for the example in Table 1. The content of the articles is not shown for brevity.
Figure 9: Chapter Thirty-Ninth (‘TORTS’) from the Civil Code, which is part of the Relevant Legislation context given to candidate LLMs, for the example in Table 1. The gold cited articles are marked in bold and the articles cited by Gemini-2.5-Flash(Figure 11) are underlined.
Figure 10: Ground truth answer by the legal expert. Text spans are highlighted in colors (green for Facts, blue for Cited Articles and orange for Analysis).
Figure 11: The answer of Gemini-2.5-Flash for the example in Table 1. Citations are highlighted in color (green for Facts and blue for Cited Articles).
Figure 12: Evaluation results for Gemini’s answer by Legal Experts and the LLM-Judge (GPT-4.1-mini Span-Judge). The response is perfect according to the legal experts. The LLM-judge is more strict and gives an 8/10 total score.