### v3

# GreekBarBench: A Challenging Benchmark for Free-Text Legal Reasoning and Citations

Abstract
We introduce GreekBarBench, a benchmark that evaluates LLMs on legal questions across five different legal areas from the Greek Bar exams, requiring citations to statutory articles and case facts. To tackle the challenges of free-text evaluation, we propose a three-dimensional scoring system combined with an LLM-as-a-judge approach. We also develop a meta-evaluation benchmark to assess the correlation between LLM-judges and human expert evaluations, revealing that simple, span-based rubrics improve their alignment. Our extensive evaluation of $13$ proprietary and open-weight LLMs shows that even though the top models exhibit impressive performance, they remain susceptible to critical errors, most notably a failure to identify the correct statutory articles.
GreekBarBench: A Challenging Benchmark for Free-Text Legal Reasoning and Citations
Odysseas S. Chlapanis1,2Dimitrios Galanis2,3Nikolaos Aletras4Ion Androutsopoulos1,2 1Department of Informatics, Athens University of Economics and Business, Greece 2Archimedes, Athena Research Center, Greece 3Athena Research Center, Greece 4University of Sheffield, United Kingdom
1 Introduction
As legal AI assistants become increasingly prevalent, the need for realistic legal LLM benchmarks has never been more imperative. [^1] https://www.abajournal.com/web/article/aba-tech-report-finds-that-ai-adoption-is-growing-but-some-are-hesitant Most widely used legal Natural Language Processing (NLP) benchmarks (Chalkidis et al., 2022; Niklaus et al., 2023) focus on classification tasks, e.g., legal judgement prediction (Aletras et al., 2016), which have been criticized (Medvedeva and Mcbride, 2023; Mahari et al., 2023) for being more constrained and less representative than real-world tasks. Even more recent LLM-focused legal benchmarks (Guha et al., 2023; Fei et al., 2024; Joshi et al., 2024) do not go beyond closed-form questions (e.g., multiple-choice questions), failing to capture the true complexity of legal reasoning in practice, which involves identifying, analyzing and synthesizing relevant information to reach a conclusion. Unfortunately, most existing benchmarks with challenging legal questions and free-text responses are proprietary and thus inaccessible to the research community. [^2] https://www.vals.ai/benchmarks
Facts
$[1]$ Antonis visited his dermatologist, Ioannis, to remove facial skin tags.
$[2]$ His assistant, Penelope, took a bottle and started washing the area.
$[3]$ The bottle contained pure acetic acid (not a solution, as prescribed), which led to burns on Antonis’ face.
$[4]$ He is now seeking €2.500 for treatment costs plus €75.000 for moral damages.
Question
Which individuals are liable for the injury?
Relevant Legal Context
Civil Code 914: Anyone who unlawfully causes damage must compensate the victim.
Civil Code 922: An employer is liable for unlawful damages caused by their employee during work.
Ground Truth Answer
Ioannis is responsible vicariously for Penelope’s
actions [3][3](Civil Code 922)and Penelope is directly
liable for her negligence [3][3](Civil Code 914).
Hence, both are liable and must compensate Antonis.
Table 1: Cropped example (English translation) from GreekBarBench. The answer requires multi-hop reasoning and citing legal articles and case facts. The spans corresponding to the scoring dimensions are highlighted in color: Facts (green), Cited Articles (blue) and Analysis (orange). Important spans are marked in bold and cited facts are denoted by square brackets. The complete example is presented in Appendix E.
Another challenge is that realistic benchmarks often require costly manual evaluation by legal experts, which limits scalability (Magesh et al., 2025; Martin et al., 2024). Automatic evaluation, using the LLM-as-a-judge framework (Zheng et al., 2023), is a promising alternative; however, its reliability has not been extensively assessed in legal reasoning (Bhambhoria et al., 2024; Li et al., 2025).
To address these issues, we present the GreekBarBench, a benchmark that evaluates the reasoning capabilities of LLMs on challenging legal questions across five legal areas. The questions are taken from the Greek Bar exams and require open-ended answers with citations to statutory articles and case facts. In addition, we introduce an accompanying benchmark for LLM-judges, designed to measure how well their scores correlate with those of human experts. GreekBarBench is the only Greek dataset for legal reasoning. Our main contributions are the following:
  • GreekBarBench: a challenging legal reasoning benchmark that requires free-text answers citing case facts and statutory articles.
  • GBB-JME: an accompanying dataset with human-evaluated answers from five different LLMs, to assess the quality of candidate LLM-judges in GreekBarBench.
  • A three-dimensional scoring system and an LLM-judge framework based on span-rubrics per dimension (Facts, Cited Articles, Analysis), which aligns well with human expert evaluation.
  • A systematic evaluation of 13 frontier and open-weight LLMs on GreekBarBench, using the best LLM-judge at GBB-JME.
  • Manual error analysis to identify important weaknesses in the model responses.
All resources including code, the two benchmarks (except for a small semi-private test set) and the prompts are publicly available. [^3] Code and prompts: https://github.com/nlpaueb/greek-bar-bench
Dataset: https://huggingface.co/datasets/AUEB-NLP/greek-bar-bench
2 GreekBarBench (GBB)
2.1 Greek Bar Exam
To become licensed attorneys, Greek law graduates must pass the Greek Bar exam, which is organized by the Greek Bar Association and requires a minimum score of six out of ten. The exam evaluates candidates through practical legal questions across five key areas of law: Civil Law, Criminal Law, Commercial Law, Public Law, Lawyers’ Code.
Greece operates under a statutory legal system, where laws are derived from codified statutes rather than judicial precedents. Consequently, the Greek Bar exam is an open-book test requiring candidates to locate and cite the correct statutory articles from the provided legal documents. These documents include the Civil Code and Civil Procedure Code, the Criminal Code and Criminal Procedure Code, eight Commercial Law codes, eleven Public Law codes, the Lawyers’ Code, and the Code of Ethics for Legal Practice (see Table 3).
Past Greek Bar exam papers and suggested solutions (2015–2024) are publicly available in an annually updated PDF booklet. [^4] The booklet is available at https://www.lawspot.gr/nomika-nea/panellinios-diagonismos-ypopsifion-dikigoron-themata-exetaseon-kai-endeiktikes-3 The authors of this paper have been granted permission by the Greek Bar Association to use the questions and by the booklet’s authors to use their respective solutions.
2.2 Task
Benchmark Lang Cite Legal Judge
Articles Context Eval
LegalBench en
LexEval (5.4) zh
CaseGen zh
OAB-Bench por
LLeQA fr
LEXam en, de
GBB (Ours) el
Table 2: Comparison of popular legal benchmarks. ‘Lang’: language of dataset. ‘Cite Articles’: the legal articles must be cited. ‘Legal Context’: the necessary legal information is provided. ‘Judge-Eval’: separate benchmark to compare LLMs-as-judges (‘–’ means that the feature is not publicly available).
The design of GreekBarBench is inspired by the workflow of candidate lawyers on the exam. They first study the case facts to identify legal issues, then navigate legal codes to locate the relevant chapter and pinpoint the specific statutory article to support their arguments. To mirror this process, each instance in the benchmark is derived from an exam question and consists of (1) the facts of the incident, (2) the exam question, and (3) a collection of potentially relevant chapters of statutory articles. The desired output is the free-text answer to the question, providing an analysis with citations to the case facts and the applicable statutory articles.
Table 2 highlights the unique contributions of GreekBarBench. It is one of the few benchmarks to require citations to legal articles (and the only one to require citations to case facts). Furthermore, it is unique in providing a corpus of potentially relevant articles and including a dedicated framework for assessing LLMs-as-judges.
2.3 Dataset Statistics
We collect a total of 65 exam papers; 13 exam papers from each of the five aforementioned areas. The PDF booklet (§ 2.1) is converted to text format and further processed (§ 2.5) to prepare the dataset. Each exam paper includes 4.7 questions on average, resulting in a total of 310 questions. We keep the questions from 2024 (22 in total) as a semi-private test set, to avoid data contamination.555‘Semi-private’ means that the test set is not public, but the raw PDFs might have leaked into the pretraining data. The public test set contains 288 questions, while the semi-private set will be updated annually with two new exam papers from each legal area.
Answering the exam questions requires citing articles from 25 legal code documents, obtained from the official website of the Greek National Printing House. [^6] https://search.et.gr/el/ Detailed statistics for these documents are presented in Table 3 per legal area. Articles are cited 931 times in total, across all exam questions. The articles within each legal code document are grouped thematically into chapters. The total number of citable articles is more than 16,000. This extensive corpus is far too large to fit entirely within the context window of any LLM.
Law Areas Questions Legal Total Cited Context
Codes Articles Articles (tokens)
Civil 71 2 3,264 286 87k
Criminal 53 2 1,253 186 58k
Commercial 58 8 4,177 159 29k
Public 71 11 2,912 118 67k
Lawyers 57 2 4,476 182 66k
Total 310 25 16,082 931 62k (avg)
Table 3: Summary of dataset statistics. ‘Legal Codes’ indicates the number of distinct legal code documents in each area. ‘Cited Articles’ is the total number of citations to legal code articles. ‘Context’ denotes the average token count of the relevant legal context (chapters of legal code) provided in the input of candidate LLMs.
2.4 Relevant Legal Context
As mentioned in Section 2.1, the Greek Bar exams are open-book, allowing candidate lawyers to navigate legal code documents to identify relevant statutory articles for the presented case. Simulating this setup presents several challenges. One approach would be implementing a Retrieval-Augmented Generation (RAG) pipeline, using sparse (e.g., BM25) or dense retrievers (Karpukhin et al., 2020) to select the $k$ most ‘relevant’ articles for inclusion in the LLM’s input. However, this approach suffers from three significant limitations: a) candidate lawyers taking the exams do not have access to such retrieval tools, making direct comparisons with human performance problematic; b) retrievers are prone to errors, creating a substantial risk that even with large values of $k$, the ground truth articles might not appear among the top retrieved articles; and c) as demonstrated by Krishna et al. (2025), benchmarking RAG systems requires testing multiple configurations with varying values of $k$ and, ideally, different retriever models, complicating fast integration of new LLMs.
To better simulate the exam environment, we propose an approach that mirrors how a candidate lawyer works: using chapter-level context rather than discrete retrieved articles. For each case, we first use regular expressions to extract the ground truth legal articles cited in its questions. We then identify the legal code chapters these articles belong to and aggregate every article from those chapters into a single corpus. This corpus serves as a broad legal context for all questions related to that case, more faithfully simulating how a lawyer navigates an entire chapter. The average corpus length is 62,000 tokens per question (Table 3), which is within the capacity of most modern LLMs.
2.5 Fact Segmentation
To streamline the evaluation process, we require citations to facts in candidate answers, though this is not mandatory in the official exams. To implement this, we segment the case facts into sentences using the Segment-Any-Text neural model (Frohmann et al., 2024) and present them as a numbered list (as shown in Table 1). This structure makes it straightforward to detect any factual errors.
2.6 Three-Dimensional Scoring System
The official evaluation committee of the Greek Bar Exams grades candidate answers on a scale of 1 to 10. This process lacks explicit guidelines and relies on a holistic comparison to the ground truth, limiting its analytical depth for benchmarking LLMs.
Drawing inspiration from established legal research and evaluation practices (Clark and DeSanctis, 2013), and guided by our legal expert annotators (§ 5.2), we develop a novel three-dimensional scoring system to improve the evaluation process for the benchmark. The proposed approach assesses legal reasoning across three dimensions: the Facts, the Cited Articles, and the Analysis. Each dimension is rated on a scale of 1 to 10, and the final score is their mean. This fine-grained framework allows the detection of specific shortcomings in the abilities of LLMs. The Facts score measures understanding of case facts; the Cited Articles score evaluates the ability to identify and cite applicable legal articles; and the Analysis score evaluates the ability to construct valid legal arguments.
3 Automatic Evaluation
To address the evaluation of free-text answers without the prohibitive cost of manual annotations, we use the LLM-as-a-judge framework (Zheng et al., 2023). LLM-judges can be categorized into two primary types: (a) pairwise LLM-judges, which evaluate two candidate answers and determine which is preferred (or declare a tie), and (b) grading LLM-judges, which assign an integer score to each individual candidate answer (Zheng et al., 2023). In our work, we focus on grading LLM-judges to allow cost-effective integration of new participant LLMs without the overhead of quadratically increasing pairwise comparisons.
To improve the alignment of LLM-judges with human expert annotators, we propose novel span-based rubrics; i.e., evaluator instructions in the form of annotated spans per question. These spans will guide the LLM-Judge in what to assess in the candidate answers. However, even with these question-specific rubrics, replicating the nuanced evaluation of human experts, especially in complex tasks like legal writing, cannot be guaranteed. For this reason, we also include a framework to meta-evaluate whether LLM-judges are suitable proxies for human evaluation on GreekBarBench.
3.1 Simple LLM-Judge
As an initial approach, we designed a straightforward prompt for a simple LLM-judge. The prompt outlines the evaluation task and explicitly defines the criteria for the Facts, Cited Articles, and Analysis scores. All necessary contextual information is provided; the facts of the case, the specific legal question, the ground truth answer with the cited articles and the candidate answer to be evaluated. This context mirrors the information provided to the human annotators for the manual evaluations (§ 5.2). The required output format is clearly specified: the model must provide an explanation for each score, followed by the integer score. The complete prompt is presented in Appendix D (Fig. 4).
3.2 Span LLM-Judge
Inspired by Clark and DeSanctis (2013), who found that rubrics improve the consistency of legal writing evaluation, we developed a rubric-based annotation process for our benchmark. Our legal experts constructed these rubrics by first identifying reference spans in the ground-truth answer corresponding to the three scoring dimensions, marking each with a distinct color: Facts (green), Cited Articles (blue), and Analysis (orange). An example of these colored spans is shown in Figure 1. Within each span, the annotators then identified important subsets, i.e., the specific words or phrases crucial for a correct answer (shown in bold in Table 1).
To minimize the annotation burden, we did not assign point values to these subsets, which contrasts with the approach in prior work (Starace et al., 2025; Pires et al., 2025). Instead, the LLM-judge is instructed to assess whether a candidate answer covers the key information within the spans, with a focus on the important subsets, and to use this assessment to score each dimension. The full prompt is available in Appendix D (Fig. 5).
4 Meta Evaluation
Meta-evaluation of grading LLM-judges aims to quantify the alignment between LLM-generated scores and human expert annotations. Previous research has predominantly relied on Pearson’s or Spearman’s correlation coefficients as primary meta-metrics (Bavaresco et al., 2024; Niklaus et al., 2025), often without substantial justification. However, advancements in meta-evaluation have emerged from the machine translation domain, particularly through the WMT Metrics Shared Task (Freitag et al., 2024, 2023), where automatic evaluation frameworks have been systematically compared and refined. The task aims to identify optimal metrics for translation quality assessment by comparing system outputs against references. Recent findings demonstrate that state-of-the-art metrics are increasingly LLM-based. The task has revealed that Pearson’s correlation coefficient exhibits vulnerability to outliers, while Spearman’s $\rho$ disregards the magnitude of ranking errors, applying uniform penalties. To address these limitations, WMT has adopted Soft Pairwise Accuracy (SPA) (Thompson et al., 2024), a metric that assigns partial credit for nearly correct rankings, thereby providing an evaluation framework that better reflects the alignment of metrics with human experts.
4.1 Soft Pairwise Accuracy (SPA)
SPA measures the degree of alignment in evaluation confidence between human experts and LLM-judges (or any other automatic metric). For example, if a human expert is confident that one system (e.g., a candidate LLM from GreekBarBench) outperforms another, but the LLM-judge is uncertain, SPA penalizes the judge—even if the ranking happened to be correct. To do this, SPA approximates the confidence of each judge (human or LLM) on each pairwise comparison between systems using p-values of appropriate permutation tests (Fisher, 1935), as detailed below. We use the original implementation. [^7] https://github.com/google-research/mt-metrics-eval Formally, SPA between a metric $m$ and human experts $h$ is defined as:
$\textit{SPA}(m,h)={\binom{N}{2}}^{-1}\sum_{i=0}^{N-1}\sum_{j=i+1}^{N-1}\left(1-\left|p_{ij}^{h}-p_{ij}^{m}\right|\right)$
where $N$ is the number of systems being evaluated, $p_{ij}^{h}$ is the p-value for the hypothesis that system $i$ is better than system $j$ according to human scores, and $p_{ij}^{m}$ is the corresponding p-value according to the metric under evaluation. The term ${\binom{N}{2}}^{-1}$ normalizes the summation by the total number of systems under comparison.
SPA permutation tests:
To estimate confidence of an evaluator (either human or automatic) in a pairwise system comparison, SPA uses permutation tests to calculate the expected mean difference under the null hypothesis that the systems are of equal quality. Specifically, a number of mock systems (1,000 in our experiments, following the original paper) are constructed as follows: for each question in the benchmark, the mock system is assigned either the score of system $i$ or system $j$ at random. The p-value is then computed as the proportion of mock systems for which the differences are greater than or equal to the mean difference between systems $i$ and $j$, as scored by the evaluator.
5 Experiments
5.1 Models
Our experiments evaluate a diverse range of LLMs, comprising proprietary models (from OpenAI, Google, Anthropic) and open-weight models; Deepseek-R1 (DeepSeek-AI et al., 2025), Gemma-3 (Team et al., 2025), and Llama-Krikri-8B (Roussis et al., 2025); a model specifically pretrained for the Greek language. We accessed proprietary models and the large open-weight Deepseek-R1 through Application Programming Interfaces (APIs) provided by OpenAI, Google, and AWS. The remaining open-weight models were deployed on a cluster of eight A100 GPUs using the vLLM framework (Kwon et al., 2023). Due to limited resources, we performed three runs for each model. We used the default parameter configurations as specified by each model’s provider.
Generation prompt:
To generate responses from candidate LLMs, we designed a system and user prompt for the benchmark. The system prompt instructs the LLM to answer with citations to Greek statutory articles. The user prompt is structured to first describe the overall task, including clear instructions on the expected output format. Then it provides the numbered facts of the case, the question and the relevant legal context. The original prompt templates are available in Appendix D.
5.2 Manual Evaluation by Legal Experts
In this section we present the manual evaluations that we collected for GBB-JME (§ 5.3), our Judge Meta-Evaluation benchmark for assessing LLM-judges on GreekBarBench. We obtain ground truth evaluations (Facts, Cited Articles, Analysis scores on a scale of 1 to 10) from two expert legal annotators who are licensed Greek lawyers with practical experience. The annotators were compensated for their time and expertise. They evaluated five LLMs on 87 questions drawn from three exam sessions (2024-A, 2023-A, and 2023-B), resulting in a total of 1,305 annotated samples. The models evaluated on all three exams were Claude-3.7-Sonnet, OpenAI-o1, GPT-4o, and Gemini-2.0-Flash. For the 2024 exam, we included the open-source Llama-3.1-70B; however, due to its poor performance and generation failures on several questions, we replaced it with Deepseek-R1 for the 2023-A and 2023-B exams. Annotations were managed with the open-source platform doccano. [^8] https://doccano.prio.org
Figure 1: Manual evaluation by legal expert annotators on the semi-private test set of the 2024 exams.
The average Krippendorff’s $\alpha$ (Krippendorff, 2011) between the two annotators on the three-dimensional scores was 0.74, and the SPA was 0.85, both indicating a substantial level of inter-annotator agreement (Artstein and Poesio, 2008). For the SPA calculation, we treated one annotator’s scores as ground truth and compared the other annotator’s scores against them. This differs from Section 5.3, where SPA measures the correlation between LLM-generated scores and the aggregated scores of human annotators.
As shown in Figure 1, the top models in the 2024 exam were Claude-3.7-Sonnet (7.55) and OpenAI-o1 (7.52). While all LLMs except Llama-3.1-70B passed the exam (score > 6.0), they still performed below the average expert (7.78) and the top 5% of experts (8.87).
5.3 Judge Meta-Evaluation (GBB-JME)
We evaluate seven LLMs-as-judges on our meta-evaluation benchmark, GBB-JME, using two prompts: the Simple-Judge3.1) and the rubric-based Span-Judge3.2). To measure the consistency of the judges, we report the average and standard error over three runs.
The results, presented in Table 4, reveal a clear trend: leading models (GPT-4.1, Gemini-2.0-Flash, and GPT-4.1-mini) generally improve with the Span-Judge prompt while weaker ones seem to struggle with its complexity. Surprisingly, the top-performing model was GPT-4.1-mini (0.862), which outperformed its larger and theoretically more powerful variant, GPT-4.1 (0.837). The stronger performance from the smaller model as a judge corroborates findings from other work (Niklaus et al., 2025).
More generally, the results highlight a key distinction between a model’s performance as a judge and as a candidate, a pattern that holds across different model families. This is illustrated by the open-weight Gemma-3-27B (0.785 with Simple-Judge), which outperformed Gemini-2.0-Flash (0.778 with Span-Judge) despite the latter’s superior performance as a candidate (see §5.4, Table 5).
Due to its strong performance and cost-effectiveness, we adopted GPT-4.1-mini with the Span-Judge prompt for all subsequent evaluations, with the total cost amounting to approximately \$100. For those seeking a more accessible option, the open-weight Gemma-3-27B serves as a strong, low-cost alternative. To support future work, we are publicly releasing the GBB-JME benchmark, enabling other researchers to evaluate new LLMs-as-judges. This approach ensures that the benchmark’s value grows over time, as it can be continuously replicated and validated with increasingly capable models serving as judges.
Model Simple-Judge Span-Judge Cost
(SPA) (SPA)
GPT-4.1-mini 0.851 ± 0.008 0.862 ± 0.001 \$\$
GPT-4.1 0.835 ± 0.010 0.837 ± 0.005 \$\$\$
Gemini-2.0-F 0.745 ± 0.009 0.778 ± 0.011 \$
Gemma-3-27B 0.785 ± 0.015 0.733 ± 0.017
Krikri-8B 0.708 ± 0.004 0.638 ± 0.019
Gemini-2.0-LF 0.667 ± 0.016 0.695 ± 0.010 \$
GPT-4.1-nano 0.675 ± 0.006 0.416 ± 0.008 \$
Table 4: Comparison of LLMs-as-judges on GBB-JME, using Simple-Judge and Span-Judge. The cost per one million input tokens is denoted by: \$ (less than \$0.3), \$\$ (less than \$1), and \$\$\$ (less than \$3).
5.4 Results on GreekBarBench
We conduct an extensive automatic evaluation of $13$ LLMs on GreekBarBench (Figure 5). We use GPT-4.1-mini as the judge with the Span-Judge prompt (§ 3.2), reporting the Facts score, Articles score, Analysis score and Mean score with standard error from three runs. The evaluation includes proprietary models such as GPT-4o, the GPT-4.1 family (GPT-4.1, GPT-4.1-mini, GPT-4.1-nano), Gemini-2.0-Flash, and Claude-3.7-Sonnet (reasoning disabled), along with the reasoning models OpenAI-o1 and Gemini-2.5-Flash. The open-weight models include the Gemma-3 family (27B, 12B, and 4B parameters), the specialized Greek model Llama-Krikri-8B-Instruct (Krikri-8B), and the reasoning model DeepSeek-R1.
Model Facts Articles Analysis Mean
Gemini-2.5-F 8.62 8.16 8.36 8.38 ± 0.02
GPT-4.1 8.65 8.05 8.34 8.35 ± 0.02
OpenAI-o1 8.24 7.55 7.55 7.78 ± 0.01
Claude-3.7-S. 8.34 7.37 7.60 7.77 ± 0.04
GPT-4.1-mini 8.28 7.18 7.40 7.62 ± 0.01
Gemini-2.0-F 8.28 7.13 7.12 7.51 ± 0.03
GPT-4o 8.15 7.10 7.18 7.48 ± 0.03
Deepseek-R1 7.65 6.48 6.55 6.90 ± 0.02
Gemma-3-27B 7.46 5.63 5.79 6.29 ± 0.02
Krikri-8B 7.20 5.65 5.74 6.20 ± 0.03
Gemma-3-12B 7.25 5.37 5.62 6.08 ± 0.04
PASS score 6.00 6.00 6.00 6.00 ± 0.00
GPT-4.1-nano 6.88 4.69 4.79 5.45 ± 0.03
Gemma-3-4B 5.53 3.76 3.95 4.42 ± 0.00
Table 5: Comparison of closed and open-weight LLMs on GreekBarBench with GPT-4.1-mini Span-Judge on three runs. The scores reported are ‘Facts’, ‘Articles’, ‘Analysis’, and ‘Mean’ score with standard error. The best scores are shown in bold, while failed scores are highlighted in red.
Gemini-2.5-Flash (8.38) and GPT-4.1 (8.35) achieved the highest scores on GreekBarBench (Table 5). These scores are not directly comparable to the human expert scores (average: 7.78, top-5%: 8.87) in Figure 1, as the models were evaluated by an LLM-judge on a larger dataset (310 vs. 22 questions). The smallest models, GPT-4.1-nano (5.45) and Gemma-3-4B (4.42) are the only models that fail the exams (passing score: 6.00). The open-weight Krikri-8B model (6.20) surpasses Gemma-3-12B (6.08) and achieves performance comparable to the significantly larger Gemma-3-27B (6.29), highlighting the benefit of language-specific pretraining.
Analyzing the scoring dimensions provides valuable insights into model capabilities. A key finding is that all models except for Gemini-2.5-Flash and GPT-4.1 struggle most with the ‘Articles‘ and the ‘Analysis‘ dimensions and that’s what separates them from the others.
Table 6 presents a fine-grained comparison for six LLMs, reporting their mean score across five different legal areas (Civil, Criminal, Commercial, Public, Lawyers) and their average (Avg). The results show that LLMs exhibit consistent performance across all legal areas. Notably, the smaller open models, Gemma-3-27B and Krikri-8B, struggle in certain areas, failing to meet the passing grade threshold of 6.00 (indicated by red). The second-best model, GPT-4.1, matches the top performer, Gemini-2.5-Flash, in ‘Criminal’ and ‘Commercial Law’, but Gemini-2.5-Flash achieves slightly higher scores in the remaining three areas.
Model Civil Crim Comm Public Lawyer Avg
Gem-2.5 8.53 8.28 8.27 8.28 8.62 8.38
GPT-4.1 8.44 8.28 8.27 8.14 8.48 8.35
Claude-3.7-S 7.72 7.37 7.31 7.79 8.29 7.77
GPT-4.1-mini 7.57 7.01 7.76 7.75 7.98 7.62
Gem-27B 6.39 5.51 5.88 6.68 7.01 6.29
Krikri-8B 5.95 5.84 5.79 6.78 6.74 6.20
Table 6: Comparison of model performance on different legal areas: ‘civil’, ‘criminal’, ‘commercial’, ‘public’, ‘lawyer’ and the average score (‘Avg’).
5.5 Chapter-based Legal Context
Model no context chapter context oracle context
f c a avg f c a avg f c a avg
Gemini-2.5-Flash 8.2 6.1 6.8 7.02 8.6 8.2 8.4 8.38 8.9 8.8 8.5 8.73
GPT-4.1 7.8 5.8 6.4 6.63 8.7 8.1 8.3 8.35 8.7 8.9 8.6 8.73
GPT-4.1-mini 7.5 4.4 5.4 5.73 8.3 7.2 7.4 7.62 8.7 8.5 8.1 8.44
Gemini-2.0-Flash 7.7 5.0 5.5 6.05 8.3 7.1 7.1 7.51 8.6 8.2 7.8 8.19
Gemma-3-27B 7.1 4.1 4.9 5.37 7.5 5.6 5.8 6.29 8.5 8.0 7.7 8.08
Table 7: Comparison with different context settings. Detailed scores for facts (f), cited articles (c), analysis (a) and mean/average (avg) are provided. The no-context setting makes the task much more challenging (most models barely get a passing score, which is above 6.00). The oracle context setting allows for higher scores, because the models make fewer mistakes in the cited articles score.
To investigate the impact of legal context, we experimented with three settings, with results presented in Table 7. The “no-context” setting provided no legal corpus, forcing models to rely solely on their parametric knowledge. The “chapter” setting, our primary setup, provided relevant articles alongside distractors, as described in Section 2.4. The “oracle” setting provided only the exact correct articles. In the no-context setting, all models except Gemini-2.5-Flash performed poorly, underscoring their limited internal knowledge of Greek legislation. Conversely, the oracle setting enabled every model to achieve excellent scores, establishing that access to the correct legal statutes is the most critical factor for success in this task.
As expected, providing more accurate legal context—progressing from the no-context to the chapters and oracle settings—dramatically improved the cited articles score for all models. The facts and analysis scores also increased slightly, showing that access to the right articles improves the model’s reasoning.
Notably, Gemini-2.0’s initial lead over GPT-4.1-mini in the no-context setting (6.05 vs. 5.73) was overturned in the chapter (7.62 vs. 7.51) and oracle (8.19 vs 8.44) settings. This divergence demonstrates that strong parametric knowledge and the ability to process long context are separate skills, as shown in (Liu et al., 2024). This highlights the necessity of evaluating both abilities independently.
6 Manual Error Analysis
To supplement our quantitative metrics, we conducted a manual error analysis to identify common failure modes. We analyzed the 10 worst-scoring responses from five key models. Human experts, guided by the scores and explanations of the LLM-judge, categorized the errors to reveal the most significant weaknesses. We focused on critical errors that led to an incorrect answer. The types of errors are the following:
Facts:
Errors include hallucinations (‘Hall.’), inventing facts not present in the provided context, and inaccuracies (‘Inac.’), misinterpreting or omitting key facts from the context.
Cited articles:
Errors include inapplicable articles (‘Inap.’), citing legal articles not relevant to the case, and missing articles (‘Missing’), failing to identify essential articles for the analysis.
Analysis:
Errors include an invalid argument (‘Argum.’), a flawed argument from misunderstood principles or fallacious reasoning, and a generic answer (‘Generic’), providing a general statement instead of an analysis tailored to the case facts.
Our analysis, while based on a small sample, highlights three primary areas of weakness: citing legal articles, understanding case facts, and constructing valid arguments. The most frequent errors related to failing to cite the correct legal articles (23). This supports our finding (§5.5) that correct article retrieval is essential. Factual (20) and analysis (17) errors were also prevalent across all models. Although rare, critical fact hallucinations (3) appeared even in top-performing models, highlighting the risks of deploying these systems without human oversight.
Model Facts Articles Analysis
Hall.
Inac.
Inap.
Missed
Argum.
Generic
Gemini-2.5-F 1 3 1 0 3 1
GPT-4.1 1 2 3 5 2 1
Claude-3.7-S 0 3 0 3 5 0
GPT-4.1-mini 0 4 1 3 4 0
Gemma-3-27B 1 5 0 7 1 0
Total 20 23 17
Table 8: Manual analysis of error types across different models. Errors are grouped into three main categories: Facts, Articles, and Analysis; and two subcategories for each one: hallucinations (‘Hall.’) and inaccuracies (‘Inac.’) for facts, inapplicable (‘Inap.’) and missed (‘Missed’) articles, and invalid argument (‘Argum.’) and generic answer (‘Generic’) for the analysis.
A key differentiator between human and model performance was the inability of LLMs to grasp subtle, legally significant context. For instance, all LLMs failed to infer that the phrase “they had differences” could indicate a motive for premeditation, which was obvious to human experts. This failure to interpret nuanced language was a common weakness, showing that significant improvement is needed for reliable use in real-world applications.
7 Related Work
Legal domain:
In the legal domain, LexGLUE (Chalkidis et al., 2022) and LEXTREME (Niklaus et al., 2023) are established benchmarks for legal classification tasks. LegalBench (Guha et al., 2023) is the standard for evaluating LLMs on legal reasoning via multiple-choice questions. More closely related to our work, task 5.4 of LexEval (Li et al., 2024) is based on Chinese Bar exams, but, unlike our approach, it does not provide citations or use LLM-as-a-judge, instead evaluating with the less reliable, overlap-based ROUGE metric (Cohan and Goharian, 2016). Similarly, LLeQA (Louis et al., 2024) contains everyday legal questions in French, but evaluation is based on the METEOR metric without measuring its correlation with human experts. CaseGen (Li et al., 2025) assesses document drafting and legal judgment generation in Chinese using the LLM-as-a-judge approach.
Two concurrent works, OAB-Bench (Pires et al., 2025) and LEXam (Fan et al., 2025), are most similar to GreekBarBench as they also provide open-ended legal questions and rubrics for LLM-judges. However, both have notable limitations. OAB-Bench, based on the Brazilian Bar exams, uses official guidelines as rubrics, but their complexity necessitates a costly model (OpenAI-o1), resulting in an evaluation cost of approximately \$50 per LLM. LEXam is based on English and German law school questions, which, unlike Bar exams, are targeted at less experienced students rather than graduates. Crucially, while both projects evaluate the quality of LLM-judges, neither provides these manual evaluations as a publicly available, standalone benchmark, preventing the community from using them to assess new judges in the future.
LLM-as-a-judge:
LLM-as-a-judge was introduced by Zheng et al. (2023), who meta-evaluated its performance against human preferences for multi-turn chat assistant dialogues. A comprehensive overview of LLM-as-a-judge and meta-evaluation resources can be found in the survey by Gu et al. (2024). Taking this concept further, JudgeBench (Bavaresco et al., 2024) introduced a general-purpose benchmark specifically for the meta-evaluation of LLM-judges. In line with our approach, other studies similarly develop separate benchmarks to meta-evaluate judges on specific tasks (Starace et al., 2025; Niklaus et al., 2025).
Evaluation Rubrics:
Legal research has for long focused on creating rubrics for consistent (human) evaluation of legal writing (Clark and DeSanctis, 2013). The Brazilian Bar exams have made their rubrics for human evaluation available, so the aforementioned OAB-Bench (Pires et al., 2025) provides them to their LLM-judges. Their rubrics consist of a manually annotated ground truth answer with comments and a table with score distributions for each element of the answer. A proprietary benchmark, BigLawBench [^9] https://www.harvey.ai/blog/introducing-biglaw-bench, describes a scoring system that uses two dimensions: the ‘source’ and ‘answer’ scores, which are analogous to our Cited Articles and Analysis. They rely on detailed instructions per question that specify explicitly the attributes that would contribute positively and negatively to the final score of candidate answers. Constructing from scratch either of these approaches is prohibitively expensive, in contrast to our simple, span-based rubrics that only require minimal annotation effort.
Greek NLP:
Important Natural Language Processing resources for the Greek language include classification models (Koutsikakis et al., 2020; Saketos et al., 2024), alongside more recent LLMs pretrained on Greek like Meltemi [^10] Meltemi was excluded from our experiments, because of its relatively small context length of 8 billion tokens. (Voukoutis et al., 2024) and Llama-Krikri-8B (Roussis et al., 2025), which we tested in our experiments (§ 5.4). Existing Greek legal datasets cover only classification and summarization tasks (Angelidis et al., 2018; Papaloukas et al., 2021; Koniaris et al., 2023). Although Greek LLM benchmarks exist for other domains, such as finance (Peng et al., 2025) and medicine (Voukoutis et al., 2024), the legal domain currently lacks one.
8 Conclusions
In this work, we introduced GreekBarBench, a benchmark evaluating LLMs on legal questions requiring citations to statutory articles and case facts. To enable robust and scalable assessment, we developed a comprehensive evaluation framework based on LLM-as-a-judge, which we validated against human experts using our GBB-JME meta-evaluation benchmark. Our results show that custom span-based rubrics significantly improve judge alignment. The extensive evaluation of 13 LLMs on GreekBarBench revealed that Gemini-2.5-Flash and GPT-4.1 achieved the best performance. Our primary finding is that success on this task is critically dependent on the quality of the provided legal context and the ability to identify the correct statutory articles within it.
Limitations
Our benchmark, GreekBarBench, assumes the availability of the relevant legal code chapters for the Relevant Legal Context component (§ 2.4). We did not evaluate the performance of retrieval models on this task, which is a critical step in real-world legal applications and could pose a significant challenge not addressed by our current setup.
A notable limitation is the cost associated with evaluating models using our framework due to the primary LLM-judge being a proprietary model (GPT-4.1-mini). To mitigate this cost, we suggest utilizing Simple-Judge with the open-weight model Gemma-3-27B. While no currently available open-weight model achieves meta-evaluation performance (SPA scores on GBB-JME) on par with GPT-4.1-mini, our public release of the benchmark and meta-evaluation dataset will allow future research to test and use more accurate and cost-effective LLM-judges.
Ethical Considerations
The development and application of legal NLP benchmarks carry significant ethical implications and potential societal impact, particularly concerning fairness, access to justice, and responsible automation (Tsarapatsanis and Aletras, 2021). Therefore, careful consideration of their design and potential uses is essential.
Our research contributes to the development of tools that could potentially assist various types of users, including legal professionals (such as judges and lawyers), students, and individuals seeking to understand legal concepts. It is crucial to emphasize that performance on this benchmark, or any similar research benchmark, should never be considered sufficient justification for deploying automated systems that substitute human experts. We strongly caution against the uncritical reliance on models evaluated solely on benchmark performance for automating legal tasks, making legal decisions, or providing legal advice.
Despite our efforts to make GreekBarBench realistic, as a research benchmark, it overlooks two critical aspects for the safe and reliable deployment of legal AI applications in practice:
  • Data Realism: Real-world legal problems are far more complex and nuanced than the structured, often simplified scenarios found in exam questions (Medvedeva and Mcbride, 2023). They often demand significant legal interpretation, ethical judgment and persuasion, particularly when the law does not provide an explicit answer for a given situation.
  • Safety: Real-world applications must ensure that the AI system handles adversarial attacks effectively. Issues like guiding the decisions of the LLMs with malicious prompting (e.g., jailbreaking), and providing confident, incorrect information when asked legally unanswerable queries are unacceptable (see discussions on AI safety principles  [^11] https://www.anthropic.com/news/core-views-on-ai-safety).
Furthermore, the primary ethical purpose of this work is not to provide a system ready for deployment, but to advance the state of legal NLP evaluation itself. By developing a benchmark that requires free-text generation, incorporates a multi-dimensional scoring system, and uses LLM-judges with explicit evaluation criteria, we aim to encourage the development of more transparent and explainable legal AI models. These features provide greater insight into how models arrive at their answers, moving beyond simple classification or multiple-choice and offering components of explainability which are crucial for gaining trust in AI applications (Medvedeva and Mcbride, 2023).
As already mentioned (§ 2.1), the authors of the solutions of the exam papers have given approval for the public reproduction of this work, with respect to the original and strictly for academic research use. Our ground truth answers are based on the year that each exam was published. This means that if the relevant laws changed in the meantime, the solutions are no longer valid. All cases in the Greek Bar exams are fictional, created solely for educational purposes, and bear no relation to real individuals or actual legal cases.
Acknowledgments
We are grateful to our legal expert annotators, Nasia Makridou and Irene Vlachou, for their diligent work, expertise, and insightful discussions, which were invaluable to this project.
This work was partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. AWS resources were provided by the National Infrastructures for Research and Technology (GRNET), with support from the EU Recovery and Resilience Facility.
Appendix A LLMs-as-judges variants
Model GPT-4.1-mini Gemma-3-27B Delta
Span Simple
Gemini-2.5-Flash 8.38 8.53 +0.15
GPT-4.1 8.35 8.38 +0.03
OpenAI-o1 7.78 8.05 +0.27
Claude-3-7-Sonnet 7.77 8.08 +0.32
GPT-4.1-mini 7.62 7.87 +0.25
Gemini-2.0-flash 7.51 7.88 +0.37
GPT-4o 7.48 7.75 +0.27
Deepseek-R1 6.90 7.21 +0.32
Gemma-3-27b 6.29 6.84 +0.55
Krikri-8B 6.20 6.80 +0.60
Gemma-3-12b 6.08 6.56 +0.48
GPT-4.1-nano 5.45 5.77 +0.31
Gemma-3-4b 4.42 5.03 +0.61
Table 9: Comparison of scores from two LLM-as-judge configurations: our primary judge (GPT-4.1-mini with Span-Judge) versus an alternative (Gemma-3-27B with Simple-Judge). The table is sorted by the GPT-4.1-mini scores. Scores from Gemma-3-27B are highlighted in bold where they alter the ranking of a model relative to its neighbors.
Table 9 presents a comparative analysis to assess the robustness of our evaluation framework. We compare the scores from our primary judge, GPT-4.1-mini with the Span-Judge prompt, against those from an alternative setup using Gemma-3-27B with a simpler Simple-Judge prompt.
Our analysis reveals two key findings. First, the overall model rankings are highly consistent across both judges, as evidenced by a strong SPA agreement of 0.92. This indicates that our top-level conclusions about model performance are robust to the choice of the judge.
Second, despite the high rank correlation, we observe a significant and systematic scoring bias. Gemma-3-27B acts as a more lenient judge, assigning scores that are, on average, 0.35 points higher than GPT-4.1-mini. Crucially, this bias is not uniform across all models. As shown in the Delta column, the score inflation is most pronounced for lower-capability models (e.g., +0.61 for Gemma-3-4b and +0.60 for Krikri-8B) and is minimal for top-tier models. The exceptionally small delta for GPT-4.1 (+0.03) is particularly noteworthy. This suggests that the judges disagree on its proximity to the top-ranked Gemini-2.5-Flash. The low delta could indicate a potential " “family bias,” where the GPT-4.1-mini judge evaluates its sibling model more favorably, rating it closer to the top performer than the external Gemma judge does.
This non-uniform bias leads to minor but notable disagreements in the fine-grained ordering of similarly-performing models. For example, the Gemma judge awards a disproportionately higher score to Claude-3-7-Sonnet and Gemini-2.0-Flash, causing them to overtake their immediate competitors in the ranking, as highlighted in bold. This experiment underscores that while overall rankings are stable, the absolute scores and the precise ordering of closely-ranked models can be sensitive to the specific judge and prompting method used.
Appendix B Contamination
Model Knowledge cutoff GreekBarBench Semi-private Delta
Gemini-2.5-Flash 2025-01-01 8.40 8.04 – 0.36
GPT-4.1 2024-06-01 8.32 7.65 – 0.67
OpenAI-o1 2023-09-01 7.77 7.51 – 0.26
Claude-3.7-Sonnet 2024-09-30 7.71 7.51 – 0.20
GPT-4.1-mini 2024-06-01 7.63 7.09 – 0.54
Gemma-3-27B 2024-08-30 6.33 5.51 – 0.82
Krikri-8B 2025-02-09 6.24 5.56 – 0.68
Semi-private set cutoff: 2024-10-03
Table 10: We report the overall performance of LLMs on the original GreekBarBench and on the semi-private set. Delta is the performance difference between the two sets. The models highlighted in bold could potentially have contaminated data and still trick the test, because their cutoff date is after the semi-private set’s date of publication.
Table 10 compares the performance of several LLMs on the public GreekBarBench and its ‘semi-private’ version. A significant drop in performance on the ‘semi-private’ set, indicated by a large negative delta, would suggest contamination. The results show no clear evidence of such contamination. The observed performance drops are minor for all models (under a 1.0-point delta), suggesting the ‘semi-private’ set is simply more difficult. The largest deltas are also associated with the lowest-performing models (Gemma-3-27B and Krikri-8B), which contradicts the expected score inflation.
The exceptions to this analysis are Gemini-2.5-Flash and Krikri-8B, whose late data cutoffs allow for potential data exposure to remain undetected. However, the concern for Krikri-8B is mitigated by its very low performance, which is inconsistent with contamination, as mentioned before. Consequently, only the results for Gemini-2.5-Flash are inconclusive. Our final assessment is that most models appear uncontaminated, while the status of Gemini-2.5-Flash cannot be determined.
Appendix C Annotator Instructions
In this section we present the instructions given to the two legal expert annotators. The annotators possessed prior experience with the evaluation task, having previously taken the exams themselves. This existing expertise allowed for concise instructions. For the general evaluation of LLM-generated answers for the manual evaluation (§ 5.2), the instruction (translated to English) was simply to:
“Evaluate the candidate answers on each scoring dimension (Facts, Cited Articles, and Analysis).”
For the creation of text spans for Span-Judge (§ 3.2), annotators were instructed to:
“Highlight the text-spans that correspond to each scoring dimension (Facts, Cited Articles, and Analysis). Highlight the most important subsets of these spans with the label important.”
Appendix D Complete Prompts
In this section we present the complete system (Fig.2) and user (Fig.3) prompts given to candidate LLMs for generation of answers, as well as the complete system prompts given to LLM-judges for the Simple- (Fig.4) and Span-Judge (Fig.5).
Figure 2: System prompt for generation given to candidate LLMs.
Figure 3: User prompt for generation given to candidate LLMs.
Figure 4: System prompt given to Simple-Judge LLM-judges.
Figure 5: System prompt given to Span-Judge LLM-judges.
Appendix E Complete Dataset Example
In this section we present the complete version of the example that we presented in Table 1. We show the complete Facts and Question (Fig. 6), the Relevant Legal Context (Figures 7 and 8), the complete Ground Truth Answer (Fig. 9), the candidate answer by Gemini-2.5-Flash (Fig. 10) and evaluations of the candidate answer by the legal experts and the LLM-judge (Fig.11).
Figure 6: Complete Facts and Question (original and below translated in English), as given in to the candidate LLMs, for the example in Table 1.
Figure 7: The Chapters of the Relevant Legislation context given to candidate LLMs, for the example in Table 1. The content of the articles is not shown for brevity.
Figure 8: Chapter Thirty-Ninth (‘TORTS’) from the Civil Code, which is part of the Relevant Legislation context given to candidate LLMs, for the example in Table 1. The gold cited articles are marked in bold and the articles cited by Gemini-2.5-Flash(Figure 10) are underlined.
Figure 9: Ground truth answer by the legal expert. Text spans are highlighted in colors (green for Facts, blue for Cited Articles and orange for Analysis).
Figure 10: The answer of Gemini-2.5-Flash for the example in Table 1. Citations are highlighted in color (green for Facts and blue for Cited Articles).
Figure 11: Evaluation results for Gemini’s answer by Legal Experts and the LLM-Judge (GPT-4.1-mini Span-Judge). The response is perfect according to the legal experts. The LLM-judge is more strict and gives an 8/10 total score.